OpenStudy (anonymous):

the sum of the measures of two acute angle is 120.the supplement of the smaller angle is 10 more than six times the complement of the larger angle.find the measures of the two acute angles.

OpenStudy (anonymous):

x < y < 90 x + y = 120 180-x = 10+ 6(90 - y)

OpenStudy (anonymous):

Add the two equations: 180 = 130 + 540 - 6y 6y = 490 y = 81 2/3 degrees x = 38 1/3 degrees

OpenStudy (anonymous):

Um...he has posted it agian.

OpenStudy (anonymous):

abtrehearn isnt it 7y=490? therefore y=70 and x=50

OpenStudy (anonymous):

Yoou're right. I fotgoy the y on the left side.